ebjohnson1:

If f(x) = 2x - 5, then f(4) is

loveboy:

f(4)=2(4)-5=3

richkid132:

Answer: you must to replace x be 4. like this f(4)=2(4)-5 so 2×4 = 8 and then 8-5=3 so the final answer is f(4)=3